William purchased x pillows at $25 each and a certain number of bolsters. Altogether, he paid $240. How many bolsters did he buy if he paid $24 for each bolster? Express your answer in termms of x.
Answer by mananth(16946)

let him buy y bolsters
pillow = $25 x numbers
bolster =$24 y numbers
25x+24y=240
add-25x
25x-25x+24y=240-25x
24y=240-25x
/24
y = 240/24 -25/24
y=10-(25x/24)